A 70 kg diver steps off of a 12 m tower and hits the water below. He travels 2.2 m below the surface before coming to a stop. Find the average force of resistance produced by the water?

We have not learned anything about PE and KE yet, only linear, projectile, and force math, how is this doable with only those?
Answer #1

let v velocity of deiver before touching the water

now apply, v^2 - u^2 = 2*a*s

v^2 - 0^2 = 2*9.8*12

v = sqrt(2*9.8*12)

= 15.34 m/s

inside the water
let a is the acceleration of the diver inside the water.

apply, vf^2 - vi^2 = 2*a*s

0^2 - 15.34^2 = 2*a*2.2

a = -15.34^2/(2*2.22)

= -53 m/s^2


so, Force exerted by the water on diver, F = m*|a|

= 70*53

= 3710 N